40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 11 02:00:21 2013, Jenne, Update, Locking, TRY gone??? 
    Reply  Sun Apr 14 15:03:17 2013, Manasa, Update, Locking, Fixing - progress 
       Reply  Sun Apr 14 17:30:14 2013, Manasa, Update, Locking, Fixed 
          Reply  Sun Apr 14 23:20:42 2013, Den, Update, Locking, Fixed 
             Reply  Thu Apr 18 21:12:56 2013, Manasa, Update, Locking, Fixed 
Message ID: 8463     Entry time: Thu Apr 18 21:12:56 2013     In reply to: 8455
Author: Manasa 
Type: Update 
Category: Locking 
Subject: Fixed 

[Den, Manasa]

TRY & TRX power measurement was redone.

TRY measures 66uW and 0.8counts on dataviewer.
TRX measures 70.4uW and 0.84counts on dataviewer.

___________________________
Detector       Power
-------------------------------------------------
QPD-Y          33uW (50%)
TRY-PD         29.8uW (49%)
Y-Camera         1%
QPD-X         35.2uW (50%)
TRX-PD        25.1uW (90%)
X-Camera    10%
____________________________

ELOG V3.1.3-